Evaluate the integral of (sin x + cosx)² dx


1
Expert's answer
2021-03-30T13:47:18-0400

"\int (\sin x + \cos x)^2\,dx = \\
\int \sin^2 x + 2\sin x\cos x + \cos^2 x\,dx = \\
\int (\sin^2 x + \cos^2 x) \,dx + \int 2\sin x \cos x \,dx = \\
\int dx + \int 2\sin x \,d(\sin x) = \\
x + \sin^2 x + C"